OpenStudy (anonymous):

I will give you a medal and be your fan!! Solve for W 1/4+W=6 and 1/3 A. w=6 1/12 B. w=6 1/7 C. w=6 2/7 D. w=6 7/12 I think it's A

OpenStudy (anonymous):

W=\[\frac{ 61 }{ 3 }-\frac{ 1 }{ 4 }\]

OpenStudy (anonymous):

60/12? Im confused

OpenStudy (anonymous):

how to get 60/12 ?

OpenStudy (anonymous):

No, i'm saying I didn't understand what you posted.. Where did you get 61 from? It's suppose to be 6 and 1/3 Not 61

OpenStudy (anonymous):

ohh sorry \[\frac{ 1 }{ 4 }+w=6*\frac{ 1 }{ 3 }\] right?

OpenStudy (anonymous):

No there's no multiplication..Im sorry that I didn't do a good job posting the question

OpenStudy (anonymous):

Do you think it's D? Im starting to think it may be D

OpenStudy (anonymous):

@esamalaa and @harrisonwalsh are you guys there?

OpenStudy (wwhitlock):

http://wolfr.am/1dVY6oH the answer is A

OpenStudy (anonymous):

ok thanks!! I went to that website to but I really wanted to learn how to actually do the problem! thanks anyways

OpenStudy (wwhitlock):

\[1/4+W=6+ 1/3\] \[1/4+W=19/3\] \[W = 19/3 - 1/4\] \[W= 76/12 - 3/12\] \[W=73/12\]